OpenStudy (anonymous):

Suppose that the graph of f contains the point (4,5). Find a point that must be on the graph of g(x)=2+f(x)

jimthompson5910 (jim_thompson5910):

Hint: if (4,5) is on the graph, then f(4) = 5

OpenStudy (anonymous):

still need help?

OpenStudy (anonymous):

yes please!

OpenStudy (anonymous):

do you understand jim thompsons' hint?

jimthompson5910 (jim_thompson5910):

Try finding g(4)

OpenStudy (anonymous):

No, I am confused about it.

OpenStudy (anonymous):

since g(x)=2+f(x), then g(4) = 2 + f(4) = ??

OpenStudy (anonymous):

Oh...shoot. My apologies I wrote my the equation wrong. It's supposed to be g(x)=2+f(6x)

OpenStudy (anonymous):

jim??? all your bud... :)

OpenStudy (anonymous):

*yours

jimthompson5910 (jim_thompson5910):

the only thing you know is that f(4) = 5 so you're stuck with f(4)...but when x = 4, f(6x) = f(6*4) = f(24)...which is in uncharted territory and we have no idea what it is (if it even exists) but you can force f(6x) to be f(4) like so 6x = 4 x = 4/6 x = 2/3 So because x = 2/3 satisfies 6x = 4, we know for a fact that f(6x) = f(4) when x = 2/3 but we also know that f(4) = 5 so this means g(x) = 2 + f(6x) g(2/3) = 2 + f(6*2/3) g(2/3) = 2 + f(4) g(2/3) = 2 + 5 g(2/3) = 7 So the point (2/3, 7) must be on the graph of g(x) = 2 + f(6x) given that (4,5) is on the graph of f(x)

OpenStudy (anonymous):

Thank you so much!